# Prominence-Aware Artifact Detection and Dataset for Image Super-Resolution

### v1

@@ Abstract
Generative image super-resolution (SR) is rapidly advancing in visual quality and detail restoration. As the capacity of SR models expands, however, so does their tendency to produce artifacts: incorrect, visually disturbing details that reduce perceived quality. Crucially, their perceptual impact varies: some artifacts are barely noticeable while others strongly degrade the image. We argue that artifacts should be characterized by their prominence to human observers rather than treated as uniform binary defects. Motivated by this, we present a novel dataset of 1302 artifact examples from 11 contemporary image-SR methods, where each artifact is paired with a crowdsourced prominence score. Building on this dataset, we train a lightweight regressor that produces spatial prominence heatmaps and outperforms existing methods at detecting prominent artifacts. We release the dataset and code to facilitate prominence-aware evaluation and mitigation of SR artifacts.
@@ 1 Introduction
Single-image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs and has become a cornerstone of low-level vision tasks. Recent progress with deep learning and generative adversarial networks (GANs) has greatly improved perceptual quality but also introduced a critical challenge: generation of visually unpleasant artifacts. These artifacts—usually unnatural patterns, smeared faces, and texture distortions—degrade the perceived quality of SISR outputs and hinder adoption. Even the latest advancements, such as transformer- and diffusion-based methods (Liang et al., 2021; Yu et al., 2024), remain prone to creating artifacts.
Despite SISR’s growing popularity, research on detecting SR artifacts remains scarce. LDL (Liang et al., 2022) and DeSRA (Xie et al., 2023a) both rely on residual statistics to localize artifacts but differ in supervision: LDL uses HR references and regularizes SR models during training, whereas DeSRA contrasts outputs from the same backbone trained with GAN vs. MSE losses and enables few-shot fine-tuning of existing SR models. Approaches such as PAL4VST (Zhang et al., 2023) cast artifact detection as a segmentation problem by predicting a binary mask from the output image.
@@
    We present the first dataset of its kind, containing 1302 SISR artifact examples with crowdsourced prominence annotations; it features diverse distortions from 11 SISR methods. To adapt for visual inspection the tight masks that artifact-detection methods produce, we propose a simple mask-postprocessing algorithm. We additionally collected prominence annotations for the 593 artifact examples in the DeSRA dataset (Xie et al., 2023a).
@@
    We propose a perceptual-prominence-modeling method that detects and quantifies artifacts in SR images and that outperforms existing approaches, judging by thorough objective and subjective evaluation. We show that our method, relative to prior art, is more applicable to fine-tuning of SR models for artifact reduction. Furthermore, we demonstrate that real-time SR can be a useful pseudo-GT for full-reference metrics.
@@
    Using our proposed artifact-detection method and existing alternatives, along with our crowdsourced prominence annotation methodology, we analyzed 11 SR methods for their propensity to generate artifacts. We show that even the newest high-quality methods such as SUPIR (Yu et al., 2024) are highly susceptible to this problem.
We published our code and datasets with prominence annotations at tinyurl.com/2u9zxtyh.
@@ 2 Related work
Recent SISR developments span Transformer-based architectures, diffusion-driven models, and hybrid generative priors. Transformer approaches such as HAT (Chen et al., 2023) and DRCT (Hsu et al., 2024b) use multiscale attention and dense-residual connections to enhance quality while remaining efficient. Diffusion-based methods like ResShift (Yue et al., 2023) and SinSR (Wang et al., 2024b) accelerate sampling via residual shifting and deterministic distillation, while StableSR (Wang et al., 2024a) leverages Stable Diffusion (Rombach et al., 2021) priors for real-world upscaling and SUPIR (Yu et al., 2024) scales models and data for text-guided photo-realistic restoration. However, these advances bring challenges, particularly the introduction of visual artifacts.
SISR evaluation has traditionally relied on full-reference metrics such as PSNR and SSIM (Wang et al., 2004), which assess reconstruction fidelity but correlate poorly with perceptual quality—especially for GAN-based outputs where details and artifacts are entangled. No-reference and perceptual metrics such as NRQM (Ma et al., 2017), LPIPS (Zhang et al., 2018), and DISTS (Ding et al., 2022) were introduced to better align with human perception and are now widely adopted in SR benchmarks. Some techniques also aim to make metrics more resistant to artifacts; for example, ERQA (Kirillova. et al., 2022) evaluates detail restoration by matching edges in reference and test images. However, in practice existing quality metrics are not sensitive enough to artifacts.
Detection and mitigation of SISR artifacts has garnered increasing attention because these artifacts reduce perceptual quality. LDL (Liang et al., 2022) predicts pixel-level artifact maps from local residual statistics. Xie et al. (2023a; b) introduced an in-lab annotated dataset with binary SR artifact masks and, building on it, proposed DeSRA that contrasts GAN-SR and MSE-SR outputs to identify artifact-prone regions, then fine-tunes the SR model on a few samples to suppress those regions.
@@
Concurrently, Ren et al. (2025) propose Hallucination Score (HS), which queries a multimodal LLM to assign a single image-level hallucination rating for SR outputs, showing strong alignment with human judgments. The main drawback of this approach is that it lacks spatial localization, which is critical for downstream tasks such as artifact mitigation, SR model fine-tuning, and for handling cases where different regions of an image exhibit different types of artifacts.
Despite these advances, the shortage of annotated datasets that explicitly focus on SR artifacts limits most approaches. While datasets such as DeSRA provide binary artifact masks, most methods remain constrained by the absence of richer annotations (e.g., prominence levels), which limits their generalization and robustness in real-world scenarios. Our work addresses this shortfall by introducing a novel dataset annotated with artifact regions and prominence scores, along with a prominence-aware detection method that supports SR fine-tuning and reveals that even the latest models like SUPIR remain highly artifact-prone.
@@ 3 Artifact dataset
Existing datasets such as DeSRA (Xie et al., 2023a) contain only binary artifact masks, without information on how noticeable the artifacts are to viewers. To enable research on artifact prominence, we introduce a dataset of 1302 artifact examples, each annotated with both a binary mask and a prominence score derived from crowdsourced assessments.
Our dataset is based on Open Images (Kuznetsova et al., 2020), a diverse collection of about 300,000 natural images (CC BY 2.0). We randomly selected 2,101 source photos, each 768×1,024 pixels. These photos then underwent 4× bicubic downsampling followed by upsampling with 11 popular SR methods (Yu et al., 2024; Wang et al., 2021b; Chen et al., 2023; Hsu et al., 2024b; Yue et al., 2023; Wang et al., 2024b; a; Cai et al., 2019; Liang et al., 2021; Wang et al., 2021a; Wu et al., 2024), yielding 23,111 images for artifact search. We obtained the initial binary artifact masks through manual annotation, and by running existing visual-quality metrics (SSIM, DISTS, LPIPS) and artifact-detection algorithms (LDL, DeSRA, in-progress versions of our method). For each of these algorithms we selected the 100 strongest detected distortions and manually discarded images without artifacts. The remaining images underwent crowdsourced prominence annotation, resulting in 697 artifact examples. The evaluation in Section 5.3 yielded 605 more examples.
@@
Figure 5: Interface presented to the viewers during subjective data collection.
Table 1: Comparison between the DeSRA and our proposed artifact dataset.
DeSRA Ours
# SR methods 3 11
# Artifacts 593 1302
Mask source in-lab automatic
Label type binary prominence
@@ 3.1 Crowdsourced annotation setup
Every image receives a ranking by 30 different participants. We compute prominence as the proportion of votes indicating the artifact is present. Before receiving access to the main questions, participants must answer four training questions, for which the correct answers are explained, followed by four test questions with hidden correct answer. Afterward, to ensure integrity, every group of 20 questions contains 4 random control questions. All responses from participants who mistakenly answer any control question are discarded. In total, 264 participants successfully completed the annotation. Appendix A analyses the impact of the participant count on the answer variability.
@@ 3.2 Mask postprocessing
An artifact-detection method aims to output a tight mask around an artifact, since doing so is more useful for further analysis and for downstream tasks such as automatic correction. But tight masks make it harder to visually judge whether the masked area contains an artifact.
To remedy this, we apply morphological operations to the masks before showing them to participants:
@@
The example in Figure 4 shows how a tight mask makes an artifact harder to notice compared with our postprocessing result. We verified the validity of this postprocessing step by running the crowdsourced prominence annotation twice on the DeSRA dataset: once with our postprocessing and once with unmodified masks. For this comparison, we ensured that separate groups of participants conducted the annotations, all while preserving the exact question order. As a result, the average artifact prominence for the entire DeSRA dataset was 49.4% with our postprocessing and 47.7% with the original masks, indicating the postprocessing helps viewers judge an artifact’s presence.
@@ 3.3 Dataset analysis
Our dataset has a bias toward zero prominence (no or unnoticeable artifacts); the reason is that we seeded the masks with results from existing image-quality metrics poorly adapted to finding artifacts and from existing artifact-detection methods that lack the ability to differentiate between barely visible and highly visible artifacts. This bias has implications for objective evaluation, potentially skewing results, and can induce overfitting during model training—a concern we address in Section H.4.
The DeSRA dataset has a more even split, with a tendency toward middle-ground artifact prominence. Notably, even though artifacts in DeSRA underwent annotations in lab, almost half of them have a prominence below 50%—that is, most viewers fail to notice them. This result confirms that binary masks are insufficient for accurate SR-artifact evaluation.
Figure 6: Architecture of the proposed artifact prominence metric. The input image is upscaled by the target SR and by RLFN as described in Section 4.3. Then, we compute three features described in Section 4.1. Finally, we run our prominence regressor described in Section 4.2.
@@ 4 Artifact prominence metric
Our goal is to predict a spatial heatmap of artifact prominence for a super-resolved image, given the corresponding low-resolution input. In this heatmap, higher values indicate regions where artifacts are more severe (i.e., more prominent to human observers).
Figure 6 shows an outline of our proposed method. Our method aggregates three features chosen from existing visual-quality metrics and artifact-detection methods. On the basis of these features, we trained a multilayer-perceptron regressor. The resulting model finds prominent image artifacts more efficiently compared with any of these features individually and compared with other approaches.
@@ 4.1 Input features
We selected features based on their proven performance for evaluating and detecting texture distortions. So, these features estimate not only visual quality of images, but also take into account structural similarity of a high-resolution image and an upscaled image.
The first feature is the visual-image-quality metric DISTS (Ding et al., 2022), which takes into account possible texture distortions and their impact on visual perception. Since DISTS is trained on natural images, it is good at detecting unnatural degradations such as SR-produced artifacts. DISTS produces a single score for a given image, so we applied it block-wise to obtain a feature map. We computed DISTS in 16×16-pixel blocks, the metric’s minimum input size.
The second feature, which we call ssm_jup, is adapted from the small-color-artifact detector from Tsereh et al. (2024), itself based on LDL (Liang et al., 2022). It targets small-scale image distortions and was shown to be effective for finding JPEG AI compression artifacts. Unlike Tsereh et al. (2024), we use all RGB channels rather than only chromatic U and V components to capture texture distortions. As with LDL, this feature requires a reference image upscaled by a more artifact-resistant method; we chose bicubic interpolation for this reference input.
The last feature, bd_jup, is the weighted sum of LPIPS (Zhang et al., 2018) and ERQA (Kirillova. et al., 2022) applied block-wise. LPIPS measures how well the upscaled image preserves perceptual quality, and is widely used in SR evaluation. Meanwhile, ERQA is designed to assess the preservation of object details and boundaries. For LPIPS, we used 32×32-pixel blocks with stride of 16. ERQA uses 8×8 blocks with no overlap. LPIPS is weighted 3:2 compared with ERQA.
4.2 Prominence regressor
We aggregated the features described above using a shallow multilayer perceptron (MLP). This network takes as input the feature values, passes them through three fully connected layers with ReLU activations, and outputs a single prominence value. The MLP independently processes each pixel of the input-feature heatmaps, predicting an artifact’s prominence in that pixel. Since our selected features employ the pixel’s neighborhood as well as wider input-image context, our final method also considers the wider context despite the MLP’s pixel-wise application.
We also experimented with CNN- and tree-based models, but the shallow MLP achieved the best overall performance (Sections H.6 and H.7).
@@ 4.3 Adapting full-resolution metrics with real-time-SR pseudo-GT
Full-reference metrics provide more-accurate detail-restoration-quality scores for SR by employing pixel-level information from the reference image. The use of such metrics in SR creates difficulties, however, since the SR-output resolution is higher than that of the original low-resolution frame.
To employ full-reference metrics, we propose the following pseudo-GT pipeline. We applied a lightweight real-time SR method to the original low-resolution frame, thereby obtaining a pseudo-GT, and then calculated the metric for this pseudo-GT and the SR output. We noticed that real-time SR methods, such as SPAN (Hu et al., 2025) and RLFN (Kong et al., 2022), produce outputs that, despite trailing heavier SR models in visual quality, are devoid of major visual artifacts. When serving as pseudo-GT for full-reference metrics, the resulting artifact-detection performance drop is small compared with using the original HR frames, as Section 5.2 shows. This approach enables our method to serve in real-world upscaling where the high-resolution GT frames are unavailable.

### v2

@@ Abstract
Generative single-image super-resolution (SISR) is advancing rapidly, yet even state-of-the-art models produce visual artifacts: unnatural patterns and texture distortions that degrade perceived quality. These defects vary widely in perceptual impact—some are barely noticeable, while others are highly disturbing—yet existing detection methods treat them equally. We propose characterizing artifacts by their prominence to human observers rather than as uniform binary defects. We present a novel dataset of 1302 artifact examples from 11 SISR methods annotated with crowdsourced prominence scores, and provide prominence annotations for 593 existing artifacts from the DeSRA dataset, revealing that 48% of them go unnoticed by most viewers. Building on this data, we train a lightweight regressor that produces spatial prominence heatmaps. We demonstrate that our method outperforms existing detectors and effectively guides SR model fine-tuning for artifact suppression. Our dataset and code are available at tinyurl.com/2u9zxtyh.
super-resolution, artifact detection, computer vision, deep learning
[…]
@@ 1 Introduction
Single-image super-resolution (SISR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs and has become a cornerstone of low-level vision tasks. While deep learning and generative adversarial networks (GANs) have greatly improved perceptual quality, they introduce a critical challenge: generation of visually unpleasant artifacts. These artifacts—usually unnatural patterns, smeared faces, and texture distortions—degrade perceived quality and hinder adoption. Even the latest transformer- and diffusion-based methods (Liang et al., 2021; Yu et al., 2024) remain prone to generating artifacts.
Despite SISR’s growing popularity, research on detecting SR artifacts remains scarce. LDL (Liang et al., 2022) and DeSRA (Xie et al., 2023) both rely on residual statistics to localize artifacts but differ in supervision: LDL uses HR references and regularizes SR models during training, whereas DeSRA contrasts outputs from the same backbone trained with GAN vs. MSE losses. Approaches such as PAL4VST (Zhang et al., 2023) predict a binary mask from the output image, similarly to semantic segmentation.
@@
    We present the first dataset of its kind, containing 1302 SISR artifact examples with crowdsourced prominence annotations; it features diverse distortions from 11 SISR methods. To adapt for visual inspection the tight masks that artifact-detection methods produce, we propose a simple mask-postprocessing algorithm. We additionally collected prominence annotations for the 593 artifact examples in the DeSRA dataset (Xie et al., 2023).
@@
    We propose a perceptual-prominence modeling method that detects and quantifies artifacts in SR images and that outperforms existing approaches, judging by thorough objective and subjective evaluation. We show that our method, relative to prior art, is more applicable to fine-tuning of SR models for artifact reduction. Furthermore, we demonstrate that real-time SR can be a useful pseudo-GT for full-reference metrics.
@@
    Using our artifact-detection method and existing alternatives, along with our prominence annotation methodology, we analyzed 11 SR methods for their propensity to generate artifacts. We show that even the newest high-quality methods such as SUPIR (Yu et al., 2024) are very susceptible to this problem.
@@ 2 Related work
Recent SISR developments span Transformer architectures, diffusion models, and hybrid generative priors. Transformers like HAT (Chen et al., 2023) and DRCT (Hsu et al., 2024) use multiscale attention and dense residual connections for efficient quality enhancement. Diffusion methods such as ResShift (Yue et al., 2023) and SinSR (Wang et al., 2024b) accelerate sampling via residual shifting and deterministic distillation. StableSR (Wang et al., 2024a) leverages Stable Diffusion (Rombach et al., 2022) priors for upscaling, and SUPIR (Yu et al., 2024) scales models and data for text-guided photo-realistic restoration. These advances, however, introduce the challenge of visual artifacts.
SISR evaluation has traditionally relied on full-reference metrics such as PSNR and SSIM (Wang et al., 2004), which assess reconstruction fidelity but correlate poorly with perceptual quality—especially for GAN-based outputs where details and artifacts are entangled. No-reference and perceptual metrics such as NRQM (Ma et al., 2017), LPIPS (Zhang et al., 2018), and DISTS (Ding et al., 2022) better align with human perception and are now widely adopted in SR benchmarks. Some techniques aim to make metrics more artifact-resistant: ERQA (Kirillova et al., 2022) evaluates detail restoration by matching edges in reference and test images. However, in practice existing metrics still fall well short of matching human perception (Borisov et al., 2025).
Detection and mitigation of SISR artifacts has garnered increasing attention because these artifacts reduce perceptual quality. LDL (Liang et al., 2022) predicts pixel-level artifact maps from local residual statistics. Xie et al. (2023) introduced an in-lab annotated dataset with binary SR artifact masks and, building on it, proposed DeSRA that contrasts GAN-SR and MSE-SR outputs to identify artifact-prone regions, then fine-tunes the SR model on a few samples to suppress those regions.
@@
Concurrently, Ren et al. (2025) propose Hallucination Score that uses a multimodal LLM to provide an image-level hallucination rating for SR outputs, showing strong alignment with human judgments. The main drawback of this approach is that it lacks spatial localization, which is critical for downstream tasks such as artifact mitigation, SR model fine-tuning, and for handling cases where different regions of an image exhibit different types of artifacts.
Despite these advances, most approaches remain limited by the shortage of annotated datasets that explicitly focus on SR artifacts. While datasets such as DeSRA provide binary artifact masks, most methods remain constrained by the absence of richer annotations (e.g., prominence levels), which limits their generalization and robustness in real-world scenarios. Our work addresses this shortfall by introducing a novel dataset annotated with artifact regions and prominence scores, along with a prominence-aware detection method that supports SR fine-tuning and reveals that even the latest models like SUPIR remain highly artifact-prone.
@@ 3 Artifact dataset
Existing datasets such as DeSRA (Xie et al., 2023) contain only binary artifact masks, without information on how noticeable the artifacts are to viewers. To enable research on artifact prominence, we introduce a dataset of 1302 artifact examples, each annotated with both a binary mask and a prominence score derived from crowdsourced assessments.
Our dataset is based on Open Images (Kuznetsova et al., 2020), a diverse collection of about 300,000 natural images (CC BY 2.0). We randomly selected 2101 source photos, each 768×1024 pixels. These photos then underwent 4× bicubic downsampling followed by upsampling with 11 popular SR methods (Yu et al., 2024; Wang et al., 2021b; Chen et al., 2023; Hsu et al., 2024; Yue et al., 2023; Wang et al., 2024b, a; Cai et al., 2019; Liang et al., 2021; Wang et al., 2021a; Wu et al., 2024), yielding 23,111 images for artifact search. We obtained the initial binary artifact masks through manual annotation, and by running existing visual-quality metrics (SSIM, DISTS, LPIPS) and artifact-detection algorithms (LDL, DeSRA, in-progress versions of our method). For each of these algorithms we selected the 100 strongest detected distortions and manually discarded images without artifacts. The remaining images underwent crowdsourced prominence annotation, resulting in 697 artifact examples. The evaluation in Section 5.4 yielded 605 more examples.
@@
Figure 3: Viewer interface for subjective data collection.
@@ 3.1 Crowdsourced annotation setup
Every image receives a ranking by 30 different participants. We compute prominence as the proportion of votes indicating the artifact is present. Before receiving access to the main questions, participants must answer four training questions, for which the correct answers are explained, followed by four test questions with hidden correct answer. Afterward, to ensure integrity, every group of 20 questions contains 4 random control questions. All responses from participants who mistakenly answer any control question are discarded. In total, 596 participants successfully completed the annotation. Appendix B analyses the impact of the participant count on the answer variability.
@@ 3.2 Mask postprocessing
An artifact-detection method aims to output a tight mask around an artifact, since doing so is more useful for further analysis and for downstream tasks such as automatic correction. But tight masks make it harder to visually judge whether the masked area contains an artifact. To remedy this, we apply morphological operations to the masks before showing them to participants:
@@
The example in Figure 4 shows how a tight mask makes an artifact harder to notice compared with our postprocessing result. We verified the validity of this step by running the crowdsourced prominence annotation twice on the DeSRA dataset: once with our postprocessing and once with unmodified masks. For this comparison, separate groups of participants conducted the annotations, with matching question order. As a result, the mean artifact prominence for the entire DeSRA dataset was 49.4% with our postprocessing and 47.7% with the original masks, indicating the postprocessing helps viewers judge an artifact’s presence.
@@ 3.3 Dataset analysis
Our dataset has a bias toward zero prominence (no or unnoticeable artifacts); the reason is that we seeded the masks with results from existing image-quality metrics poorly adapted to finding artifacts and from existing artifact-detection methods that lack the ability to differentiate between barely visible and highly visible artifacts. This bias has implications for objective evaluation, potentially skewing results, and can induce overfitting during model training—a concern we address in Section I.4.
The DeSRA dataset shows a more balanced distribution, centered around moderate artifact prominence. Notably, despite its lab annotations, almost half of them have a prominence below 50%—that is, most viewers fail to notice them. This confirms that binary masks are insufficient for accurate SR-artifact evaluation.
Table 1: Comparison between the DeSRA and our artifact dataset.
DeSRA Ours
# SR methods 3 11
# Artifacts 593 1302
Mask source in-lab automatic
Label type binary prominence
Figure 6: Architecture of the proposed artifact prominence metric. The input image is upscaled by the target SR and by RLFN as described in Section 4.3. Then, we compute three features described in Section 4.1. Finally, we run our fusion module described in Section 4.2.
@@ 4 Artifact prominence metric
Our goal is predicting a spatial artifact prominence heatmap for a super-resolved image from its low-resolution input, where higher values indicate more severe artifacts (more prominent to human observers). Figure 6 outlines our method, which aggregates three features from existing quality and artifact-detection metrics via a lightweight fusion module. The resulting model finds prominent artifacts more efficiently than any individual feature or other approaches.
@@ 4.1 Input features
We selected features based on their proven performance for evaluating and detecting texture distortions. These features estimate not only the visual quality, but also the structural similarity between the reference and the upscaled image.
The first feature is DISTS (Ding et al., 2022), a visual-image-quality metric which accounts for texture distortions and their perceptual impact. As DISTS is trained on natural images, it effectively detects unnatural degradations like SR artifacts. DISTS produces a single image-level score, so we computed it block-wise in 16×16-pixel blocks, the minimum input size of the metric.
The second feature, which we call ssm_jup, is adapted from the small-color-artifact detector from Tsereh et al. (2024), itself based on LDL (Liang et al., 2022). It targets small-scale distortions and was shown to be effective for finding JPEG AI artifacts. To capture texture distortions, we modify the detector to use all RGB channels rather than only chromatic U and V components. Like LDL, this feature requires a reference image upscaled by an artifact-resistant method; we chose bicubic interpolation for this input.
The last feature, bd_jup, is a weighted sum of LPIPS (Zhang et al., 2018) and ERQA (Kirillova et al., 2022) applied block-wise. LPIPS measures how well the upscaled image preserves perceptual quality, and is widely used in SR evaluation. Meanwhile, ERQA assesses the preservation of object details and boundaries. For LPIPS, we used 32×32-pixel blocks with stride of 16. ERQA uses 8×8 blocks with no overlap. LPIPS is weighted 3:2 compared with ERQA.
Appendix A details the implementation of ssm_jup and bd_jup.
4.2 Fusion module
We experimented with several architectures for fusing the features described above into a single prominence prediction, including CNN-based and tree-based models (see Sections I.6 and I.7). A shallow multilayer perceptron (MLP) achieved the best overall performance, so we adopted it as our feature fusion module.
The MLP takes as input the feature values, passes them through three fully connected layers (3-128-128-1) with ReLU activations, and outputs a single prominence value. It independently processes each pixel of the input-feature heatmaps, yet still captures broader context since the features themselves encode both the pixel’s neighborhood and wider image-level information.
@@ 4.3 Adapting full-resolution metrics with real-time-SR pseudo-GT
Full-reference metrics provide more-accurate detail restoration quality scores for SR by using pixel-level information from the reference image. The use of such metrics in SR creates difficulties, however, since the SR-output resolution is higher than that of the original low-resolution frame.
To employ full-reference metrics, we propose the following pipeline. We applied a lightweight SR method to the original low-resolution frame, thereby obtaining a pseudo-GT, and then calculated the metric for this pseudo-GT and the SR output. We noticed that real-time SR methods, such as SPAN (Wan et al., 2024) and RLFN (Kong et al., 2022), produce outputs that, despite trailing heavier SR models in visual quality, are devoid of major visual artifacts. When serving as pseudo-GT for full-reference metrics, the artifact-detection performance drop is small compared with using the original HR frames, as Section 5.2 shows. This approach enables our method to serve in real-world upscaling where the high-resolution GT frames are unavailable.